# Yongjiang Wua, Yongtao Lia, Lihua Fenga, Jiuqiang Liub, Guihai Yuc, [^2] Corresponding author

### v2

@@ 1.2 Applications: Stabilities for Erdős–Ko–Rado’s theorem
Next suppose that the equality holds. Then $\mathcal{F}(\bar{n})=\left\{F_{1},F_{2},F_{3}\right\}$ and for any permutation $\sigma$ on $[3]$, we have $|\mathcal{G}_{\sigma(2)}|=\binom{n-k-t-2}{k-2}$ and $|\mathcal{G}_{\sigma(3)}|=\binom{n-k-t-3}{k-3}$. If there exist $i\neq j\in[3]$ such that $\left|F_{i}\cap F_{j}\right|\leq k+t-2$, then let $\sigma=(1i)(2j)$. So there exist $y_{1}\neq y_{2}\in\mathcal{F}_{\sigma(1)}$ such that $y_{1},y_{2}\notin F_{\sigma(2)}$. This implies that $|\mathcal{G}_{\sigma(2)}|\geq\binom{n-k-t-2}{k-2}+\binom{n-k-t-3}{k-2}$, a contradiction. Thus $\mathcal{F}(\bar{n})=\left\{F_{1},F_{2},F_{3}\right\}$ with $\left|F_{i}\cap F_{j}\right|=k+t-1$ for any $1\leq i<j\leq 3$. We claim that $|\cap_{i=1}^{3}F_{i}|=k+t-1$. Otherwise we can write $F_{1}=M\cup\{a,b\},F_{2}=M\cup\{a,c\},F_{3}=M\cup\{b,c\}$, where $|M|=k+t-2$ and $a,b,c$ are pairwise different. But then for any permutation $\sigma$ on $[3]$, we have $|\mathcal{G}_{\sigma(3)}|\geq\binom{n-k-t-2}{k-2}$, a contradiction. Therefore, by the shiftedness of $\mathcal{F}$, we conclude that the equality holds if and only if $\mathcal{F}(\bar{n})=\left\{F_{1},F_{2},F_{3}\right\}$ with $F_{i}:=[k+t-1]\cup\{k+t-1+i\}$ for each $i\in[3]$ and $\mathcal{G}(n)=\left\{G\in\binom{[n-1]}{k-1}:G\cap F_{i}\neq\emptyset\text{ for any }i=1,2,3\right\}$. It follows that
@@
By Theorem 1.2 and the shiftedness of $\mathcal{F}$, the equality holds if and only if $\mathcal{F}(n)=\left\{[k+t-1]\right\}$ and $\mathcal{G}(n)=\left\{G^{*}\in\binom{[n-1]}{k-1}:G^{*}\cap[k+t-1]\neq\emptyset\right\}$. Note that $\mathcal{F}(\bar{n})$ and $\mathcal{G}(n)$ are cross-intersecting. By the shiftedness of $\mathcal{F}$, if $\mathcal{F}(n)=\left\{[k+t-1]\right\}$ and $\mathcal{G}(n)=\left\{G^{*}\in\binom{[n-1]}{k-1}:G^{*}\cap[k+t-1]\neq\emptyset\right\}$, then $\mathcal{F}(\bar{n})=\left\{[k+t-1]\cup\{i\}:i\in\{k+t,\ldots,n-1\}\right\}$, $\mathcal{G}(\bar{n})=\left\{G\in\binom{[n-1]}{k}:G\cap[k+t-1]\neq\emptyset\right\}$.
@@
$\leq\binom{n}{k}-\binom{n-k-t}{k}-\binom{n-k-t-1}{k-1}-\binom{n-k-t-2}{k-2}+3.$
If the equality holds, then $1=\binom{n-k-t-3}{k-4}$, i.e., $k=4$, and =$\mathcal{F}=\{[t+3]\cup\{i\}:i\in\{t+4,\ldots,n\}\}$ and $\mathcal{G}=\left\{G\in\binom{[n]}{4}:G\cap[t+3]\neq\emptyset\right\}$.
@@
3 Revisiting stabilities of intersecting families
As promised, we will present a unified framework to revisit several stability results for the Erdős–Ko–Rado theorem. To begin with, we fix the following notations.
@@ 3.1 Short proof of the Hilton–Milner theorem
Case (B). Let $\mathcal{G}\subseteq\binom{[n]}{k}$ be obtained from $\mathcal{F}$ by applying some shifting operations such that $\mathcal{G}$ is not EKR, but after the shifting operation $s_{i,j}$, the resulting family $s_{i,j}(\mathcal{G})$ is EKR. In this case, all sets of $s_{i,j}(G)$ contain the element $i$, it follows that $\mathcal{G}(\bar{i},\bar{j})=\emptyset$ and $\{i,j\}\cap G\neq\emptyset$ for every $G\in\mathcal{G}$. By the maximality of $|\mathcal{G}|$, we have $\mathcal{G}(i,j)=\binom{[n]\backslash\{i,j\}}{k-2}$. Note that $\mathcal{G}(i,\bar{j})\subseteq\binom{[n]\backslash\{i,j\}}{k-1}$ and $\mathcal{G}(\bar{i},j)\subseteq\binom{[n]\backslash\{i,j\}}{k-1}$ are non-empty cross-intersecting. So Theorem 1.1 can be applied. Since $\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})=\emptyset$, the inequality in Theorem 1.1 holds strictly. Then
@@ 3.2 Short proof of the Han–Kohayakawa theorem
To make the proof clearer, we divide it into case $k\geq 4$ and case $k=3$. Firstly, let us consider the case $k\geq 4$ and present some tools as follows.
@@
Then $\mathcal{F}$ and $\mathcal{G}\subseteq\binom{[n]}{\ell}$ are cross-intersecting if and only if $\mathcal{F}\cap\mathcal{D}_{k}(\mathcal{G})=\emptyset$ or equivalently $\mathcal{G}\cap\mathcal{D}_{\ell}(\mathcal{F})=\emptyset$. Moreover, for given $\mathcal{F}$, $\mathcal{G}=\binom{[n]}{\ell}\backslash\mathcal{D}_{\ell}(\mathcal{F})$ is the largest family that is cross-intersecting with $\mathcal{F}$, and for given $\mathcal{G}$, $\mathcal{F}=\binom{[n]}{k}\backslash\mathcal{D}_{k}(\mathcal{G})$ is the largest family that is cross-intersecting with $\mathcal{G}$.
@@
Moreover, for $n>2k$, the equality holds if and only if =$\mathcal{F}^{\prime}=\left\{[k]\cup\{i\}:i\in\{k,k+1\}\right\}$ and $\mathcal{G}^{\prime}=\left\{G\in\binom{[n]}{k}:G\cap([k]\cup\{i\})\neq\emptyset,i\in\{k,k+1\}\right\}$, or two more possibilities when $k=3$, namely, =$\mathcal{F}^{\prime}=\{[2]\cup\{i\}:i\in\{3,4,\ldots,n\}\}\text{ and }\mathcal{G}^{\prime}=\left\{G\in\binom{[n]}{3}:G\cap[2]\neq\emptyset\right\}$, or $\mathcal{F}^{\prime}=\mathcal{G}^{\prime}=\left\{G\in\binom{[n]}{3}:1\in G\right\}$. We need to return to the structure of $\mathcal{F}$ and $\mathcal{G}$. In the first extremal case, we have $|\mathcal{F}|=|\mathcal{F}^{\prime}|=2$. Denote =$\mathcal{F}=\left\{F_{1},F_{2}\right\}$. Then $\left|F_{1}\cap F_{2}\right|=k-1$ otherwise the upper bound would be smaller. Thus $\mathcal{G}=\left\{G\in\binom{[n]}{k}:G\cap F_{i}\neq\emptyset,i\in[2]\right\}$. In the second extremal case, we have $|\mathcal{F}|=|\mathcal{F}^{\prime}|=n-2={n-2\choose 3-2}$ and $|\mathcal{G}|=|\mathcal{G}^{\prime}|={n\choose 3}-{n-2\choose 3}$. For a fixed family $\mathcal{F}\subseteq\binom{[n]}{3}$, the maximality of $|\mathcal{F}|+|\mathcal{G}|$ yields $\mathcal{G}=\binom{[n]}{3}\setminus\mathcal{D}_{3}(\mathcal{F})$. So we get $|\mathcal{D}_{3}(\mathcal{F})|={n-2\choose 3}$, which implies that the equality case of Lemma 3.1 occurs. Thus, it follows from Lemma 3.1 that =$\mathcal{F}=\{[2]\cup\{i\}:i\in\{3,4,\ldots,n\}\}$ and $\mathcal{G}=\left\{G\in\binom{[n]}{3}:G\cap[2]\neq\emptyset\right\}$ under isomorphism. In the third extremal case, we have $|\mathcal{F}|=|\mathcal{G}|={n-1\choose 2}$. Similarly, we get $|\mathcal{D}_{3}(\mathcal{F})|=\binom{n-1}{3}$. By Lemma 3.1, we obtain =$\mathcal{F}=\left\{F\in\binom{[n]}{3}:1\in F\right\}$ and $\mathcal{G}=\left\{G\in\binom{[n]}{3}:1\in G\right\}$ under isomorphism. ∎
@@
We first have that $s_{i,j}(\mathcal{G})$ is HM at $i$. Then $|\mathcal{G}(\bar{i},\bar{j})|\leq 1$. By Case 2, we may assume that $s_{i,j}(\mathcal{G})$ is not EKR. If $|\mathcal{G}(\bar{i},\bar{j})|=0$, then $|\mathcal{G}(i,\bar{j})|\geq 2$ and $|\mathcal{G}(\bar{i},j)|\geq 2$. In addition, we have $|\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})|=1$. Observe that the inequality in Lemma 3.2 still holds strictly for $\mathcal{G}(i,\bar{j})$ and $\mathcal{G}(\bar{i},j)$. So the same argument with Case 2 works. Now assume that $|\mathcal{G}(\bar{i},\bar{j})|=1$. Since $\mathcal{G}(i,j)$ and $\mathcal{G}(\bar{i},\bar{j})$ are cross-intersecting, we have $|\mathcal{G}(i,j)|\leq\binom{n-2}{k-2}-\binom{n-k-2}{k-2}.$ In addition, by the properties of $\mathcal{G}$, $\mathcal{G}(i,\bar{j})$ and $\mathcal{G}(\bar{i},j)$ are cross-intersecting, $\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})=\emptyset$, $|\mathcal{G}(i,\bar{j})|\geq 1$ and $|\mathcal{G}(\bar{i},j)|\geq 1$. The inequality in Theorem 1.1 holds strictly. Then $|\mathcal{G}(i,\bar{j})|+|\mathcal{G}(\bar{i},j)|\leq\binom{n-2}{k-1}-\binom{n-k-1}{k-1}$. Consequently,
@@
In order to prove Theorem 1.6 for $k=3$, we need to establish the corresponding versions of $k=2$ for Theorem 1.3 and Lemma 3.2. For simplicity, we postpone the proof in Appendix B.
@@ 3.3 Short proof of the Huang–Peng theorem
Note that for $32k+1\leq n\leq 3k-$ 3, $\left|\mathcal{K}_{2}(n,k)\right|\geq\left|\mathcal{J}_{3}(n,k)\right|$, and the equality holds only for $k=4$; for $n\geq 3k-2$, $\left|\mathcal{K}_{2}(n,k)\right|<\left|\mathcal{J}_{3}(n,k)\right|$. Therefore, in the folllowing, we may assume that for any $x\in[n]$, there are at least three sets in $\mathcal{F}$ not containing $x$. Then it suffices to show that
@@
$|\mathcal{F}|=|\mathcal{F}(\bar{1})|+|\mathcal{F}(1)|\leq\binom{n-1}{k-1}-\binom{n-k-1}{k-1}-\binom{n-k-2}{k-2}-\binom{n-k-3}{k-3}+3,$
and for $n\geq 2k+1$, the equality holds if and only if $\mathcal{F}\cong\mathcal{J}_{3}(n,k)$ or when $k=5$, $\mathcal{F}\cong\mathcal{G}_{4}(n,5)$.
@@
By conditions, we have $\mathcal{G}(\bar{i},\bar{j})\subseteq\{E_{1},E_{2}\}\text{ for some }|E_{1}\cap E_{2}|=k-1$ and so $|\mathcal{G}(\bar{i},\bar{j})|\leq 2$. By Cases 2 and 3, we may assume that $s_{i,j}(\mathcal{G})$ is neither EKR nor HM. Therefore, if $|\mathcal{G}(\bar{i},\bar{j})|=0$, then $|\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})|=2$. In addition, $|\mathcal{G}(i,\bar{j})|\geq 3$ and $|\mathcal{G}(\bar{i},j)|\geq 3$. By Lemma 3.3, the same argument with Case 2 works. If $|\mathcal{G}(\bar{i},\bar{j})|=1$, then $|\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})|=1$. In addition, we have $|\mathcal{G}(i,\bar{j})|\geq 2$ and $|\mathcal{G}(\bar{i},j)|\geq 2$. By Lemma 3.2, the same argument with Case 3 works.
@@
Note that $\mathcal{G}_{1}\cap\mathcal{G}_{2}=\emptyset$, $\mathcal{G}(i,j)$ and $\mathcal{G}(\bar{i},\bar{j})$ are cross-intersecting. Then we have
@@
It remains to prove Theorem 1.7 for $k=4$. We shall establish the corresponding versions of Theorem 1.4 and Lemma 3.3 for $k=3$. We defer the detailed discussions to Appendix C.

### v3

@@ 1.2 Applications: Stability for Erdős–Ko–Rado’s theorem
Next suppose that the equality holds. Then $\mathcal{F}(\bar{n})=\left\{F_{1},F_{2},F_{3}\right\}$ and for any permutation $\sigma$ on $[3]$, we have $|\mathcal{G}_{\sigma(2)}|=\binom{n-k-t-2}{k-2}$ and $|\mathcal{G}_{\sigma(3)}|=\binom{n-k-t-3}{k-3}$. If there exist $i\neq j\in[3]$ such that $\left|F_{i}\cap F_{j}\right|\leq k+t-2$, then let $\sigma=(1i)(2j)$. So there exist $y_{1}\neq y_{2}\in\mathcal{F}_{\sigma(1)}$ such that $y_{1},y_{2}\notin F_{\sigma(2)}$. This implies that $\left\{G\in\binom{[n-1]}{k-1}:y_{1}\in G\text{ or }y_{2}\in G,G\cap F_{\sigma(2)}=\emptyset\right\}\subseteq\mathcal{G}_{\sigma(2)}$. Thus $|\mathcal{G}_{\sigma(2)}|\geq\binom{n-k-t-2}{k-2}+\binom{n-k-t-3}{k-2}$, a contradiction. Thus $\mathcal{F}(\bar{n})=\left\{F_{1},F_{2},F_{3}\right\}$ with $|F_{1}\cap F_{2}\cap F_{3}|=k+t-1$. Otherwise we can write $F_{1}=M\cup\{a,b\},F_{2}=M\cup\{a,c\},F_{3}=M\cup\{b,c\}$, where $|M|=k+t-2$ and $a,b,c$ are pairwise different. By symmetry, we may assume that $F_{i}=F_{\sigma(i)}$ for $i\in[3]$. Then $\left\{G\in\binom{[n-1]}{k-1}:a\in G,G\cap F_{\sigma(3)}=\emptyset\right\}\subseteq\mathcal{G}_{\sigma(3)}$. Hence, for any permutation $\sigma$ on $[3]$, we have $|\mathcal{G}_{\sigma(3)}|\geq\binom{n-k-t-2}{k-2}$, a contradiction. Therefore, by the shiftedness of $\mathcal{F}$, we conclude that the equality holds if and only if $\mathcal{F}(\bar{n})=\left\{F_{1},F_{2},F_{3}\right\}$ with $F_{i}:=[k+t-1]\cup\{k+t-1+i\}$ for each $i\in[3]$ and $\mathcal{G}(n)=\left\{G\in\binom{[n-1]}{k-1}:G\cap F_{i}\neq\emptyset\text{ for any }i=1,2,3\right\}$. It follows that
@@
Since $k-1\geq 3$, by Theorem 1.2, the equality holds if and only if $\mathcal{F}(n)=\left\{F_{1}\right\}$ for some $F_{1}\in\binom{[n-1]}{k+t-1}$ and $\mathcal{G}=\left\{G\in\binom{[n-1]}{k-1}:G\cap F_{1}\neq\emptyset\right\}$. Then the shiftedness of $\mathcal{F}$ leads to $\mathcal{F}(n)=\left\{[k+t-1]\right\}$ and $\mathcal{G}(n)=\left\{G^{*}\in\binom{[n-1]}{k-1}:G^{*}\cap[k+t-1]\neq\emptyset\right\}$. Note that $\mathcal{F}(\bar{n})$ and $\mathcal{G}(n)$ are cross-intersecting. By the shiftedness of $\mathcal{F}$, if $\mathcal{F}(n)=\left\{[k+t-1]\right\}$ and $\mathcal{G}(n)=\left\{G^{*}\in\binom{[n-1]}{k-1}:G^{*}\cap[k+t-1]\neq\emptyset\right\}$, then $\mathcal{F}(\bar{n})=\left\{[k+t-1]\cup\{i\}:i\in\{k+t,\ldots,n-1\}\right\}$, $\mathcal{G}(\bar{n})=\left\{G\in\binom{[n-1]}{k}:G\cap[k+t-1]\neq\emptyset\right\}$.
@@
$\leq\binom{n}{k}-\binom{n-k-t}{k}-\binom{n-k-t-1}{k-1}-\binom{n-k-t-2}{k-2}+3,$
where the second equality holds by applying the Pascal identity five times. Moreover, if all the equalities above hold, then $1=\binom{n-k-t-3}{k-4}$, i.e., $k=4$, and =$\mathcal{F}=\{[t+3]\cup\{i\}:i\in\{t+4,\ldots,n\}\}$ and $\mathcal{G}=\left\{G\in\binom{[n]}{4}:G\cap[t+3]\neq\emptyset\right\}$.
@@
3 Revisiting stability of intersecting families
As promised, we will present a unified framework to revisit several stability results for the Erdős–Ko–Rado theorem. To begin with, we fix the following notation.
@@ 3.1 Short proof of the Hilton–Milner theorem
Case (B). Let $\mathcal{G}\subseteq\binom{[n]}{k}$ be obtained from $\mathcal{F}$ by applying some shifting operations such that $\mathcal{G}$ is not EKR, but after the shifting operation $s_{i,j}$, the resulting family $s_{i,j}(\mathcal{G})$ is EKR. In this case, all sets of $s_{i,j}(\mathcal{G})$ contain the element $i$, it follows that $\mathcal{G}(\bar{i},\bar{j})=\emptyset$ and hence $\{i,j\}\cap G\neq\emptyset$ for every $G\in\mathcal{G}$. Moreover, $\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})=\emptyset$ holds. By the maximality of $|\mathcal{G}|$, we have $\mathcal{G}(i,j)=\binom{[n]\backslash\{i,j\}}{k-2}$. Since $\mathcal{G}$ is not EKR, we get that $\mathcal{G}(i,\bar{j})$ and $\mathcal{G}(\bar{i},j)$ are non-empty. Note that $\mathcal{G}(i,\bar{j})\subseteq\binom{[n]\backslash\{i,j\}}{k-1}$ and $\mathcal{G}(\bar{i},j)\subseteq\binom{[n]\backslash\{i,j\}}{k-1}$ are cross-intersecting. So Theorem 1.1 can be applied. Since $\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})=\emptyset$, the inequality in Theorem 1.1 holds strictly. Then
@@
Remark. After submitting this paper, we are informed by Bulavka and Woodroofe [3] that their recent work on cross-intersecting families also yields a short proof of the Hilton–Milner theorem. We remark here that the main result [3, Theorem 2] is a direct consequence of Theorem 1.2. To some extent, our framework used in the above proof surprisingly develops a unified approach for deriving deep stability results for intersecting families, not just the Hilton–Milner theorem. For instance, we will present more potential applications of our approach in next two subsections.
@@ 3.2 Short proof of the Han–Kohayakawa theorem
To make the proof clearer, we divide it into case $k\geq 4$ and case $k=3$ (See Appendix A). Here we consider the case $k\geq 4$ and present some tools. Note that the case when $k=3$ is in the appendix by something along the lines of this subsection.
@@
Then $\mathcal{F}$ and $\mathcal{G}\subseteq\binom{[n]}{\ell}$ are cross-intersecting if and only if $\mathcal{F}\cap\mathcal{D}_{k}(\mathcal{G})=\emptyset$ or equivalently $\mathcal{G}\cap\mathcal{D}_{\ell}(\mathcal{F})=\emptyset$. Moreover, for given $\mathcal{F}$, $\mathcal{G}=\binom{[n]}{\ell}\backslash\mathcal{D}_{\ell}(\mathcal{F})$ is the largest $\ell$-uniform family that is cross-intersecting with $\mathcal{F}$, and for given $\mathcal{G}$, $\mathcal{F}=\binom{[n]}{k}\backslash\mathcal{D}_{k}(\mathcal{G})$ is the largest $k$-uniform family that is cross-intersecting with $\mathcal{G}$.
@@
Moreover, for $n>2k$, the equality holds if and only if =$\mathcal{F}^{\prime}=\left\{[k]\cup\{i\}:i\in\{k,k+1\}\right\}$ and $\mathcal{G}^{\prime}=\left\{G\in\binom{[n]}{k}:G\cap([k]\cup\{i\})\neq\emptyset,i\in\{k,k+1\}\right\}$, or two more possibilities when $k=3$, namely, =$\mathcal{F}^{\prime}=\{[2]\cup\{i\}:i\in\{3,4,\ldots,n\}\}\text{ and }\mathcal{G}^{\prime}=\left\{G\in\binom{[n]}{3}:G\cap[2]\neq\emptyset\right\}$, or $\mathcal{F}^{\prime}=\mathcal{G}^{\prime}=\left\{G\in\binom{[n]}{3}:1\in G\right\}$. We need to return to the structure of $\mathcal{F}$ and $\mathcal{G}$. In the first extremal case, we have $|\mathcal{F}|=|\mathcal{F}^{\prime}|=2$. Denote =$\mathcal{F}=\left\{F_{1},F_{2}\right\}$. Then $\left|F_{1}\cap F_{2}\right|=k-1$ by the case of equality in Theorem 1.3. Thus $\mathcal{G}=\left\{G\in\binom{[n]}{k}:G\cap F_{i}\neq\emptyset,i\in[2]\right\}$. In the second extremal case, we have $|\mathcal{F}|=|\mathcal{F}^{\prime}|=n-2={n-2\choose 3-2}$ and $|\mathcal{G}|=|\mathcal{G}^{\prime}|={n\choose 3}-{n-2\choose 3}$. For a fixed family $\mathcal{F}\subseteq\binom{[n]}{3}$, the maximality of $|\mathcal{F}|+|\mathcal{G}|$ yields $\mathcal{G}=\binom{[n]}{3}\setminus\mathcal{D}_{3}(\mathcal{F})$. So we get $|\mathcal{D}_{3}(\mathcal{F})|={n-2\choose 3}$, which implies that the equality case of Lemma 3.1 occurs. Thus, it follows from Lemma 3.1 that =$\mathcal{F}=\{[2]\cup\{i\}:i\in\{3,4,\ldots,n\}\}$ and $\mathcal{G}=\left\{G\in\binom{[n]}{3}:G\cap[2]\neq\emptyset\right\}$ under isomorphism. In the third extremal case, we have $|\mathcal{F}|=|\mathcal{G}|={n-1\choose 2}$. Similarly, we get $|\mathcal{D}_{3}(\mathcal{F})|=\binom{n-1}{3}$. By Lemma 3.1, we obtain =$\mathcal{F}=\left\{F\in\binom{[n]}{3}:1\in F\right\}$ and $\mathcal{G}=\left\{G\in\binom{[n]}{3}:1\in G\right\}$ under isomorphism. ∎
@@
We first have that $s_{i,j}(\mathcal{G})$ is HM at $i$. Then $s_{i,j}(\mathcal{G})\subseteq\big{\{}F\in\binom{[n]}{k}:i\in F,F\cap B\neq\emptyset\big{\}}\cup\{B\}$ for some $i\notin B$, This implies that $\mathcal{G}(\bar{i},\bar{j})=s_{i,j}(\mathcal{G}(\bar{i},\bar{j}))\subseteq\{B\}$ and hence $|\mathcal{G}(\bar{i},\bar{j})|\leq 1$. By Case 2, we may assume that $s_{i,j}(\mathcal{G})$ is not EKR. If $|\mathcal{G}(\bar{i},\bar{j})|=0$, then $|\mathcal{G}(i,\bar{j})|\geq 2$ and $|\mathcal{G}(\bar{i},j)|\geq 2$. In addition, we have $|\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})|=1$. Observe that the inequality in Lemma 3.2 still holds strictly for $\mathcal{G}(i,\bar{j})$ and $\mathcal{G}(\bar{i},j)$, and $|\mathcal{G}(i,j)|\leq\binom{n-2}{k-2}$. So the same argument with Case 2 works. Now assume that $|\mathcal{G}(\bar{i},\bar{j})|=1$. Since $\mathcal{G}(i,j)$ and $\mathcal{G}(\bar{i},\bar{j})$ are cross-intersecting, we have $|\mathcal{G}(i,j)|\leq\binom{n-2}{k-2}-\binom{n-k-2}{k-2}.$ In addition, by the properties of $\mathcal{G}$, $\mathcal{G}(i,\bar{j})$ and $\mathcal{G}(\bar{i},j)$ are cross-intersecting, $\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})=\emptyset$, $|\mathcal{G}(i,\bar{j})|\geq 1$ and $|\mathcal{G}(\bar{i},j)|\geq 1$. The inequality in Theorem 1.1 holds strictly. Then $|\mathcal{G}(i,\bar{j})|+|\mathcal{G}(\bar{i},j)|\leq\binom{n-2}{k-1}-\binom{n-k-1}{k-1}$. Consequently,
@@
In order to prove Theorem 1.6 for $k=3$, we need to establish the corresponding versions of $k=2$ for Theorem 1.3 and Lemma 3.2. For simplicity, we postpone the proof in Appendix A.
@@ 3.3 Short proof of the Huang–Peng theorem
Then $\left|\mathcal{K}_{2}(n,k)\right|\geq\left|\mathcal{J}_{3}(n,k)\right|$ if and only if $\binom{n-1}{k-1}-2\binom{n-k-1}{k-1}+\binom{n-k-3}{k-1}+2\geq\binom{n-1}{k-1}-\binom{n-k-1}{k-1}-\binom{n-k-2}{k-2}-\binom{n-k-3}{k-3}+3$, and if and only if $\binom{n-k-3}{k-3}\geq\binom{n-k-3}{k-2}+1$, and if and only if $32k+1\leq n\leq 3k-$ 3. Moreover, the equality holds only for $k=4$ and $n=9$. In addition, for $n\geq 3k-2$, $\left|\mathcal{K}_{2}(n,k)\right|<\left|\mathcal{J}_{3}(n,k)\right|$. Therefore, in the following, we may assume that for any $x\in[n]$, there are at least three sets in $\mathcal{F}$ not containing $x$. Then it suffices to show that
@@
$|\mathcal{F}|=|\mathcal{F}(\bar{1})|+|\mathcal{F}(1)|\leq\binom{n-1}{k-1}-\binom{n-k-1}{k-1}-\binom{n-k-2}{k-2}-\binom{n-k-3}{k-3}+3.$
Moreover, for $n\geq 2k+1$, the equality holds if and only if
$\mathcal{F}(\overline{1})$ =$=\left\{\{2,\ldots,k,k+1\},\{2,\ldots,k,k+2\},\{2,\ldots,k,k+3\}\right\},$
$\mathcal{F}(1)$ =$=\left\{G\in\binom{[n]\setminus\{1\}}{k-1}:G\cap([2,k]\cup\{i\})\neq\emptyset\text{ for }i=k+1,k+2,k+3\right\},$
or one more possibility when $k-1=4$, i.e., $k=5$,
$\mathcal{F}(\overline{1})=\left\{[2,5]\cup\{i\}:i\in\{6,\ldots,n\}\right\},\ \mathcal{F}(1)=\left\{G\in\binom{[n]\setminus\{1\}}{4}:G\cap[2,5]\neq\emptyset\right\}.$
It follows that
$\mathcal{F}=$$\left\{G\in\binom{[n]}{k}:1\in G,G\cap([2,k]\cup\{i\})\neq\emptyset\text{ for }i=k+1,k+2,k+3\right\}$
$\cup\{2,\ldots,k,k+1\}\cup\{2,\ldots,k,k+2\}\cup\{2,\ldots,k,k+3\},$
or one more possibility when $k=5$,
=$\mathcal{F}=\left\{G\in\binom{[n]}{5}:1\notin G,[2,5]\subseteq G\right\}\cup\left\{G\in\binom{[n]}{5}:1\in G,G\cap[2,5]\neq\emptyset\right\}.$
For the former, let $x_{0}=1$, $E=[2,k]$ and $J_{3}=\{k+1,k+2,k+3\}$. Then $\mathcal{F}=\mathcal{J}_{3}(n,k)$. For the latter, let $x_{0}=1$ and $E=[2,5]$. Then $\mathcal{F}=\mathcal{G}_{4}(n,5)$. Moreover, for any $\mathcal{F}^{\prime}\subseteq\binom{[n]}{k}$ with $s_{ij}(\mathcal{F}^{\prime})=\mathcal{F}$, we have $\mathcal{F}^{\prime}\cong\mathcal{J}_{3}(n,k)$ or one more possibility when $k=5$, $\mathcal{F}^{\prime}\cong\mathcal{G}_{4}(n,5)$.
@@
Since $s_{i,j}(\mathcal{G})\subseteq\mathcal{J}_{2}(n,k)$, we have $\mathcal{G}(\bar{i},\bar{j})\subseteq\{E_{1},E_{2}\}\text{ for some }|E_{1}\cap E_{2}|=k-1$ and so $|\mathcal{G}(\bar{i},\bar{j})|\leq 2$. By Cases 2 and 3, we may assume that $s_{i,j}(\mathcal{G})$ is neither EKR nor HM. Therefore, if $|\mathcal{G}(\bar{i},\bar{j})|=0$, then $|\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})|=2$. In addition, $|\mathcal{G}(i,\bar{j})|\geq 3$ and $|\mathcal{G}(\bar{i},j)|\geq 3$. By Lemma 3.3, the same argument with Case 2 works. If $|\mathcal{G}(\bar{i},\bar{j})|=1$, then $|\mathcal{G}(\bar{i},j)\cap\mathcal{G}(i,\bar{j})|=1$. In addition, we have $|\mathcal{G}(i,\bar{j})|\geq 2$ and $|\mathcal{G}(\bar{i},j)|\geq 2$. By Lemma 3.2, the same argument with Case 3 works.
@@
Clearly, $|\mathcal{G}_{1}|=\binom{n-k-2}{k-2}$ and $|\mathcal{G}_{2}|=\binom{n-k-3}{k-3}$. Note that $\mathcal{G}_{1}\cap\mathcal{G}_{2}=\emptyset$, $\mathcal{G}(i,j)$ and $\mathcal{G}(\bar{i},\bar{j})$ are cross-intersecting. Then we have
@@
It remains to prove Theorem 1.7 for $k=4$. We shall establish the corresponding versions of Theorem 1.4 and Lemma 3.3 for $k=3$. We defer the detailed discussions to Appendix B.
The covering number of a family is the size of the smallest set that intersects all sets from the family. There are also some literature [9, 20, 21, 36] studying the intersecting families with covering number at least three or five. Therefore, it is natural to investigate the following problem.
Problem 3.4.
What kind of inequalities on cross-intersecting families would be needed to provide new proofs of the results on intersecting families with covering number at least three?